OpenStudy (anonymous):

What is the area of a circle with the diameter of 12

OpenStudy (anonymous):

113.04

OpenStudy (saifoo.khan):

Radius = 6

OpenStudy (anonymous):

d=12 r=6 pir^2=36pi

OpenStudy (anonymous):

pi * r^2 r = radius = half of diameter so r = 6 area = pi * 6^2

OpenStudy (saifoo.khan):

\[Area = \pi r^2\]\[\pi 6^2\]\[36\pi\]

OpenStudy (turingtest):

\[A= \pi r^2\]\[r=d/2\]\[A= \pi d^2/4=\pi 12^2/4=36 \pi\]